OpenStudy (anonymous):

One of two supplementary angles is five times the other. Find the angles. (Recall that supplementary angles are angles whose sum is 180 degrees.) Which equation can be used to solve for the smallest angle? 5x = 180 6x = 180 2x + 5 = 180

OpenStudy (zpupster):

x the original angle and 5x the original angle can you see it now??

OpenStudy (anonymous):

yes i c what u r doing.... but what is the equation????

OpenStudy (zpupster):

don't you think that you should add the 2 angles to get 180

OpenStudy (anonymous):

yes, but my multiply choice doesn't gave me 5x+2=180

OpenStudy (zpupster):

x+5x = 180 or 6x = 180 take the coefficients of x and add them 1+5=6

OpenStudy (anonymous):

thankx...